OpenStudy (anonymous):

What is the sum of an 8-term geometric series if the first term is 15 and the last term is -4,199,040?

OpenStudy (anonymous):

So \[ar ^{1} = 15\] and \[ar ^{8} = -4,199,040\] Solve for r an a \[\frac{ r ^{8} }{ r} = \frac{ -4,199,040 }{ 15 }\] so solve and find that r = -6 so sub back r and find a, a= -2.5 now sub those values into \[Sn = \frac{ 1-r ^{n} }{ 1-r }\] solve for Sn which is the sum of the 8 terms.